OpenStudy (anonymous):

PTG has side lengths of 12 cm, 25 cm, and 19 cm. Which type of triangle is PTG?

OpenStudy (ryan123345):

Hey there! I see that you are new to openstudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! I wish you good luck on openstudy and dont forget to read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Hope you have fun!

OpenStudy (ryan123345):

and plug in the numbers in the pythagorean theorem. if it is true it is a tight triangle

OpenStudy (ryan123345):

if the first number is bigger than its acute and if the first number is smaller than its acute

OpenStudy (ryan123345):

lets try it

OpenStudy (ryan123345):

\[12^2+25^2=19^2\] \[144+625=361\] \[769=361\] so since the first number is bigger its a acute triangle